OpenStudy (anonymous):

Someone help me simplify an expression! Someone help me simplify an expression! @Mathematics

OpenStudy (anonymous):

\[\sqrt[4]{36}\times \sqrt[4]{9}\div \sqrt[4]{4}\]

OpenStudy (anonymous):

that expression is equal to \[^4\sqrt{\frac{36*9}{4}}\]\[=^4\sqrt{81}\]\[=3\]

OpenStudy (anonymous):

Thank you!!! :D